OpenStudy (coledrf):

i need help! divide (7t^5+14t^4-28t^3+35t^2) divided by 7t^2

OpenStudy (loser66):

suggestion: break it out: I do the first one for you as example \(\dfrac{7t^5}{7t^2}=t^3\) do the same with the 2nd term, the 3rd term,...etc then put them back as if we never break them up before, hihihihi...